OpenStudy (anonymous):

Nick follows a rule to deposit money into his savings account each month. The past 4 months, he deposited 6,10 ,14 , and 18 dollars. His sister determines that each deposit is increased by 6, so his next deposit should be 18 + 6=$24 . What error did she make? What should his next deposit be? Explain your answer.

OpenStudy (anonymous):

The question isn't typed out correctly.

OpenStudy (anonymous):

there fixed it

OpenStudy (anonymous):

What is \(10-6\)?

OpenStudy (math2400):

Isn't it increasing by 4 after 6? Cuz 10+4=14 and 14+4=18 so 18+4=22 the next deposit should be 22 because she calculated wrong (assumed that he was adding 6 every time because that was what he initially deposited)

OpenStudy (anonymous):

that is correct i just asked my dad and he read your answer math2400 andd that you looked at the problem with great perspective thank you

OpenStudy (math2400):

no problem(: glad i could help!